# Distributed and heterogeneous tensor–vector contraction algorithms for high performance computing

### v1

@@ 5.2 State-of-the-art performance of the tensor–vector contraction
We find it convenient to give the reader a notion about the expected parallel performance of the dTVC algorithm using the software available in the literature. To this end, we provide a comparison of our proposed solution against the massively parallel tensor contraction framework (CTF [23]). This framework enables hybrid parallelism via MPI and MKL libraries and also makes extensive use of the optimized kernels provided by the later, namely: BLAS, LAPACK, and ScaLAPACK. We compiled CTF against the latest Intel software suite available on MN4 (Intel oneAPI 2022.3) and the recommended high performance tensor transpose (HPTT) library for maximum performance. CTF can be regarded as a DSL for tensor algebra that allows, for instance, to effortlessly express a 2-mode tensor–vector contraction (i.e., $\mathcal{Y}=\mathcal{A}\times_{3}\mathbf{x}$) within a single line of C++ code:
@@
assuming that $\mathcal{A}$ (i.e., A["ijkl"]) is a fourth-order tensor. There are two important aspects about the previous line of code that are worth discussing. Firstly, both tensors and the vector are implicitly distributed among different MPI processes and, therefore, this operation is not exposed to the user (CTF claims to minimize communication costs). Secondly, each tensor operation such as the one indicated above is susceptible [^2] This will depend on the contraction mode, the tensor splitting assumed by CTF, and the number of MPI processes. to group assemble the resulting partial subtensors $\mathcal{Y}^{(p)}$, in order to form the global tensor $\mathcal{Y}$ (i.e., Y["ijl"]). Therefore, benchmarking such a CTF code example is similar to timing the distributed contraction in addition to the tensor assembly. This global gathering incurs MPI communication, which can consume more computational time than the tensor contraction itself. This is undesirable, especially in the case of the dHOPM3, where the partial subtensors resulting from each dTVC operation do remain distributed throughout the entire execution (see Algorithm 1).
@@
Figure 3 presents a strong scalability performance comparison between our proposed dTVC algorithm and the CTF implementation using up to 8 MPI processes (4 compute nodes) on MN4. We show the normalized throughput measured in it/sp because, contrarily to CTF, our implementation seeks one-dimensional optimal splittings (see Fig. 1); consequently, order 7–8 tensors are contracted with 7 processes, while the ninth-order tensor only exploits 5 processes. Next, for each tensor in Fig. 3, the corresponding results are the aftermath of averaging the normalized throughput values over all contraction modes and all splitting dimensions ($d^{2}$ times) in the case of our implementation. Since CTF splits the tensors internally, the associated results are only averaged over $d$ contraction modes. In both cases, the overhead related to the tensor assembly discussed above is accounted for.
@@
We shall start by focusing on the results of Fig. 3 corresponding to a single MPI process, where only shared-memory parallelism is available and there is no need to form a global output tensor. In this situation, our distributed implementation is able to sustain more than 8 it/sp across all tensors. On the other hand, CTF only remains competitive for the matrix case (order 2) and gradually loses performance, being more than four times slower for the tenth-order tensor due to the subpar performance of MKL’s batch-strided kernels employed for the internal contraction modes [8]. Getting back to the matrix case using our proposed implementation (solid bars), it can be readily seen that increasing the number of distributed processes has no visible impact on performance: indeed, these results are all above 80% of the theoretical peak performance of the machine. This is expected because the amount of memory being communicated using one-dimensional splitting is reduced by a factor of 30623×30623$\times$30623 × with respect to the input tensor size (7.5 GB). In contrast, much smaller factors are associated with higher order tensors (e.g., 8×8$\times$8 × for the tenth-order tensor) and thus larger communication overheads as it can be inferred from the remaining solid bars of Fig. 3. In this regard, CTF presents a rather strange behaviour for third- and especially second-order input tensors with throughput values well below 0.5 it/sp, thereby making it unsuitable for distributed matrix–vector (or vector–matrix) multiplication of large square matrices. On the other hand, CTF starts approaching and can even surpass the performance of our implementation for higher order tensors and more than four MPI processes. But such cases are largely dominated by the time spent assembling the global output tensor, which can be more than one order of magnitude higher than that of the TVC operation itself. Indeed, the bandwidth associated with $d>6$ and $p=8$ do not attain 10% of the peak performance and hence cannot be considered representative of HPC scenarios. Finally, it is worth mentioning that we have not optimized neither the disjoint union, Eq. (1), nor the global reduction, Eq. (2), operations involved in the formation of the global output tensor.
@@
Figure 3 emphasises the criticality of the distributed tensor–vector contraction and, particularly, its associated global reconstruction process. In this sense, the convenience of expressing tensor contractions straightforwardly through a DSL (via CTF, tiledArray and similar libraries), can be counterbalanced by the lack of performance arising from the continuous assembly of the distributed output subtensors. In the remainder of this work, we will evaluate the performance of the dTVC algorithm without reconstruction. As previously discussed, this operation is irrelevant in the dHOPM3 because it naturally exploits the distributive property of the tensor contraction.
@@ 5.3 Distributed native tensor–vector contraction performance
Table 2 reports on the results of the dTVC ignoring the global construction of the final tensor, which allow us to focus exclusively on the computational kernel performance. In particular, we are concerned about the matriziced tensor’s shape —tall and skinny vs. short and fat matrices— on the contraction performance. In addition to this, we propose the comparison of the looped approach based on state-of-the-art MKL kernels against our native implementation using two compilers. We remind the reader that results are averaged over all contraction modes and all splitting dimensions and normalized by the theoretical peak bandwidth corresponding to up to eight MPI processes on MN4 (order 7–9 tensors exploit lower MPI processes due to the one-dimensional optimal splitting, as previously discussed).
It can be readily seen on Table 2 that the looped dTVC algorithm based on MKL’s batch-strided kernel yields the worst results, with relatively low average bandwidth and high standard deviation rates. The high variability of the results, except in the matrix case, evidence the lack of mode-oblivious properties of this implementation despite the fact that it is based upon a carefully optimized kernel. As demonstrated in Ref. [6], MKL is more sensitive to the shape of the matricized tensor and, in the case of concatenated small MVMs, LIBXSMM may be able to report slightly better figures. In any case, they are not expected to remain competitive against those shown in the next two columns, which belong to our native dTVC implementation using the same compilation stack as well as the LLVM compilation infrastructure. Differences between compilers are minimum and, in both cases, our proposed algorithm achieves almost three quarters of the theoretical peak performance (1024 GB/s) with a standard deviation below 10%. We remind that this low value accounts for changes not only in the contraction mode, but also in the splitting dimension which ultimately affects the shape of the matricized form of the tensors. Therefore, we can affirm that the distributed strategy proposed in this work remains oblivious to these two variables.
Table 2: dTVC normalized, averaged bandwidth (measured as a percentage of the theoretical peak value) and corresponding unbiased, sample standard deviation percentage (within brackets) over all contraction modes and all splitting dimensions for the tensors of Table 1. Results correspond to up to 8 MPI processes on MN4 and two TVC algorithms (looped and native).
Order Looped (MKL) Native (Intel) Native (LLVM)
@@
Figure 4(a) shows that, in general, the averaged bandwidth of all tensor contractions is between 70% to 85% of the theoretical peak on MN4, which is in good agreement with the STREAM benchmark on a single NUMA node of this architecture (81.6%). This rather small performance variability of about 15% is inevitable: for instance, not all the tensors have the exact same memory footprint as reported in Table 1. With more than 16 MPI processes, the strong scalability performance begins to increase slightly as the bandwidth bottleneck moves from main memory to L3 cache. Figure 4(b) shows absolute dTVC performance values in the form of an averaged throughput measured in kernel invocations per second (it/s). We can see how the curves associated with each tensor are very close to each other except for $d=9,10$ due to their larger memory footprint (in relative terms: 6.7% and 22.7%, respectively). We confirm the linear (and superlinear) scalability already predicted by the bandwidth curves. In the light of the results gathered in Figs. 4(a)–(b), it is the first time that a distributed tensor–vector contraction remains oblivious to MPI splitting, contraction mode, and also tensor order, provided that the tensor size does not vary significantly.
@@
Figures 4(c)–(f) report on the dTVC performance for the CTE-ARM and CTE-POWER systems. We use the native, TVC algorithm to benchmark the ARM architecture while GPUs are evaluated against a looped TVC implementation making use of cuBLAS. Starting with the CPU case, Fig. 4(c) shows slightly decreasing values of the normalized bandwidth, especially for $d>3$, but they are overall within the range of 50% to 70% of the theoretical peak (on average, STREAM reports 60% on a single node). The almost linear throughput curves of Fig. 4(d) are a bit more spread than those of MN4 but, once more, we can say that our novel implementation remains oblivious to splitting, contraction mode and tensor order for this particular architecture. Finally, note that although CTE-ARM starts with better absolute throughput values over MN4 thanks to its high end memory, its overall worse strong scalability yields subpar performance with 128 MPI processes.
@@
Figure 4(e) shows the dTVC performance on GPUs using state-of-the-art cuBLAS kernels. We observe a large variability going from 90% of the theoretical bandwidth peak to only 20% (on average, STREAM reports 61.8% on a single graphics card). Interestingly, only the last two higher order tensors report somewhat constant bandwidth figures, but 10% to 15% below the reference STREAM value. This is expected since batched kernels do incur some overhead while enqueueing consecutive multiplication kernels within a unique function call. This distributed GPU implementation based on looped TVC lacks the obliviousness of its CPU counterpart implementation. In absolute terms, see Fig. 4(f), the CTE-POWER attains maximum throughput values around 4096 it/s, a quantity that is not much higher than the one reported by a general-purpose architecture endowed with DDR4 memory. It is worth highlighting that dTVC remains an embarrassingly parallel applications (the final disjoint union operation is discarded) with zero MPI communication parts and, consequently, the scalability drop in Fig. 4(f) is likely to be attributed to the rather low device occupancy. This supports our choice of one-dimensional tensor splittings since very fine-grain parallelization tends to hamper performance, especially on this type of accelerators.
(a) Full dHOPM3 execution ($i=0$)
(b) Zoom-in on the last $k$ internal contractions of iteration $j=0$
Figure 5: Paraver traces from the hybrid execution of (a) the dHOPM3 algorithm and (b) a zoom-in view of the last $k$ internal contractions of iteration $j=0$ using the tenth-order tensor of Table 1 split along its last dimension. These traces were obtained on MN4 using 4 MPI processes, 8 cores per process, and 32 OmpSs-2 tasks per process for illustration purposes.
5.4 Distributed, native higher order power method performance
@@
We take the opportunity to describe herein the shared-memory parallel strategy adopted for dHOPM3 in which every contraction (lines 7, 9, 10 and 12 of Algorithm 1) takes advantage of task-based parallelization. In contrast to ubiquitous OpenMP parallel for loops, tasks permit to overlap computations corresponding to different TVC instances, thereby maximizing the parallel performance over fork-join. This key novel parallel design is clearly seen in Fig. 5, which corresponds to a real execution carried out on MN4, instrumented with Extrae and visualized with Paraver [^3] Extrae and Paraver are available at: https://tools.bsc.es/ for a tenth-order tensor distributed among four processes. While Fig. 5(a) covers a full dHOPM3 execution ($i=0$), the figure below is a zoomed version towards the end of the first external iteration $j=0$. The overlapping of several TVC operations represented by tasks of different colors is clearly established. White gaps evidence code sections with no CPU usage and correspond exclusively to the final normalization phase (lines 13–17 of Algorithm 1) where synchronous collective communication takes place and MPI processes must wait for the slowest one due to system imbalance. Both subfigures also reveal how computational costs greatly reduce after each contraction as the resulting tensor loses one dimension and its size decreases by about one order of magnitude. On the other hand, a fork-join, shared-memory strategy and, by extension a CUDA-based implementation, presents one additional synchronization per contraction, resulting in overall lower parallelism with respect to our task-based implementation. Table 3 demonstrates how a genuine data-flow execution based on either OpenMP (OMPtk) or OmpSs-2 tasks (OSStk) can provide, on average, up to a 10% increase in memory bandwidth over a classical strategy using “parallel for” constructs (OMPfj). It is worth mentioning that these moderate speedups are limited by the synchronous MPI collectives at the end of each external iteration of dHOPM3. In cases where the HOPM requires many iterations to reach converge ($i\gg 1$), task-based relaxation techniques similar to those that have been put in place in classical linear algebra iterative algorithms [33] could be employed to remove global synchronizations and approximate the base of output vectors. Such a case is beyond the scope of this work, but represents a real scenario where task-based parallelization could yield additional advantages.
Table 3: dHOPM3 normalized bandwidth (measured as a percentage of the theoretical peak value) over the last splitting dimension for the tensors of Table 1. Results correspond to up to 8 MPI processes on MN4 and three parallel implementations.
@@
Figure 6(a) shows the normalized bandwidth of dHOPM3 on MN4, which presents similarities with the dTVC bandwidth results previously shown in Fig. 4(a). The variability range is expected to widen due to the continuous decrease of the tensor size after each contraction but, in general, normalized values are between 60% to 85% most of the time. Contrarily to dTVC, dHOPM3 is not an embarrassingly parallel application and, sometimes, the cost of synchronous communications can be as important as computations as it is the case with $d=2$ and $p>32$. When looking at the absolute throughput, see Fig. 6(b), it starts at about 4–8 it/s and reaches between 512–1024 it/s with 128 MPI processes for $d=3,4$ while it drops by a factor of two for the matrix case due to excessive synchronization. It can be readily seen that Figs. 4(b) and 6(b) are reasonably equivalent with the particularity that there exists a factor above two between them. This is because the computational time of dHOPM3 is primarily dominated by two contractions, particularly $k=1$ for $j=0$ and $k=0$ for $j=1$, as illustrated in Fig. 5(a).
@@
Figures 6(c)–(f) show the dHOPM3 results on the other two architectures equipped with high bandwidth memory. In the case of the CTE-ARM cluster, each contraction makes use of our native TVC algorithm and no significant differences can be appreciated between Figs. 4(c) and 6(c). Nevertheless, it is true that, in the case of lower order tensors, dHOPM3 bandwidth values are more sensible to the number of processes due to synchronous communication. A similar factor above 2×2$\times$2 × between dTVC and dHOPM3 throughput is found on CTE-ARM. Looking at the results obtained on these two CPU-based architectures, it can be inferred that the dHOPM3 algorithm inherits the obliviousness of the native dTVC kernel it is built upon.
@@
Figure 6(e) shows the dHOPM3 normalized bandwidth under Volta GPUs. Compared to the dTVC results of Fig. 4(e), dHOPM3 performance is better in the sense that curves do remain closer to each other, especially with moderate values of $p$. This is in contrast to what is observed on the two previous architectures. As a reminder, dTVC results were averaged over all the splitting dimensions while dHOPM3 figures are only calculated for the last splitting dimension for minimal memory footprint. For this reason, one can expect higher bandwidth values for dHOPM3 up to a certain number of processes; otherwise, the relatively low GPU occupancy combined with the increasing synchronization costs of collectives results in an important reduction in bandwidth, with values as low as 10% with 128 MPI processes. It is worth noting that these are indeed the best results one can achieve in terms of communication: device memory is transferred between devices directly via the NCCL library that provides genuine nonblocking calls to MPI functions from the CPU host. Other communication approaches, not shown here for the sake of brevity, always result in subpar performance. Finally, Fig. 6(f) confirms that the maximum throughput achieved on GPUs is above 512 it/s and does not necessarily correspond to the largest number of processes. Considering all the strong scalability results presented in this section, none of the clusters equipped with high bandwidth memory is able to surpass a general purpose supercomputer with regular DDR4 memory. This is mainly attributed to the low capacity HBM2 modules installed on these systems.
@@
Figure 7: dHOPM3 normalized throughput for the tensors of Table 1 stored in different floating-point precision formats and split along their last dimension. Results correspond to up to 8 MPI processes on (a) MN4 and (b) CTE-ARM.

### v2

@@ 5.2 State-of-the-art performance of the tensor–vector contraction
We find it convenient to give the reader a notion about the expected parallel performance of the dTVC algorithm using the software available in the literature. To this end, we provide a comparison of our proposed solution against the massively parallel tensor contraction (CTF) framework [23] (Deinsum [25] does not currently support tensor–vector contractions). This framework enables hybrid parallelism via MPI and MKL libraries and also makes extensive use of the optimized kernels provided by the later, namely: BLAS, LAPACK, and ScaLAPACK. We compiled CTF against the latest Intel software suite available on MN4 (Intel oneAPI 2022.3) and the recommended high performance tensor transpose (HPTT) library for maximum performance. CTF can be regarded as a DSL for tensor algebra that allows, for instance, to effortlessly express a 2-mode tensor–vector contraction (i.e., $\mathcal{Y}=\mathcal{A}\times_{2}\mathbf{x}$) within a single line of C++ code:
@@
assuming that $\mathcal{A}$ (i.e., A["ijkl"]) is a fourth-order tensor. There are two important aspects about the previous line of code that are worth discussing. Firstly, both tensors and the vector are implicitly distributed among different MPI processes and, therefore, this operation is not exposed to the user (CTF claims to minimize communication costs). Secondly, each tensor operation such as the one indicated above is susceptible [^5] This will depend on the contraction mode, the tensor splitting assumed by CTF, and the number of MPI processes. to assemble the resulting partial subtensors $\mathcal{Y}^{(p)}$, in order to form the global tensor $\mathcal{Y}$ (i.e., Y["ijl"]). Therefore, benchmarking such a CTF code example is similar to timing the distributed contraction in addition to the tensor assembly. This global gathering incurs MPI communication, which can consume more computational time than the tensor contraction itself. This is undesirable, especially in the case of the dHOPM3, where the partial subtensors resulting from each dTVC operation do remain distributed throughout the entire execution (see Algorithm 1).
@@
Figure 4 presents a strong scalability performance comparison between our proposed dTVC algorithm and the CTF implementation using up to 8 MPI processes (4 compute nodes) on MN4. We show the normalized throughput measured in it/sp because, contrarily to CTF, our implementation seeks one-dimensional optimal splittings (see Fig. 1). Consequently, order 7–8 tensors are contracted with 7 processes, while the ninth-order tensor only exploits 5 processes. Next, for each tensor in Fig. 4, the corresponding results are the aftermath of averaging the normalized throughput values over all contraction modes and all splitting dimensions ($d^{2}$ times) in the case of our implementation. Since CTF splits the tensors internally, the associated results are only averaged over $d$ contraction modes. In both cases, the overhead related to the tensor assembly discussed above is accounted for.
@@
We shall start by focusing on the results of Fig. 4 corresponding to a single MPI process ($p=1$), where only shared-memory parallelism is available and there is no need to form a global output tensor. In this situation, our distributed implementation is able to sustain more than 8 it/sp across all tensors. On the other hand, CTF only remains competitive for the matrix case (order 2) and gradually loses performance, being more than four times slower for the tenth-order tensor due to the subpar performance of MKL’s batch-strided kernels employed for the internal contraction modes [8]. Getting back to the matrix case using our proposed implementation (solid bars), it can be readily seen that increasing the number of distributed processes has no visible impact on performance: indeed, the corresponding results are all above 80% of the theoretical peak performance of the machine. This is expected because the amount of memory being communicated using one-dimensional splitting is reduced by a factor of 30623×30623$\times$30623 × with respect to the input tensor size (7.5 GB). In contrast, much smaller factors are associated with higher-order tensors (e.g., 8×8$\times$8 × for the tenth-order tensor) and thus larger communication overheads as it can be inferred from the remaining solid bars of Fig. 4. In this regard, CTF presents a rather strange behaviour for third- and especially second-order input tensors with throughput values well below 0.5 it/sp, thereby making it unsuitable for distributed matrix–vector (or vector–matrix) multiplication of large square matrices. On the other hand, CTF starts approaching and can even surpass the performance of our implementation for higher-order tensors and more than four MPI processes. But such cases are largely dominated by the time spent assembling the global output tensor, which can be more than one order of magnitude higher than that of the TVC operation itself. Indeed, the bandwidth associated with $d>6$ and $p=8$ do not attain 10% of the peak performance and hence cannot be considered representative of HPC scenarios. Finally, it is worth mentioning that we have not optimized neither the disjoint union, Eq. (1), nor the global reduction, Eq. (2), operations involved in the formation of the global output tensor.
@@
Figure 4 emphasises the criticality of the distributed tensor–vector contraction and, particularly, its associated global reconstruction process. In this sense, the convenience of expressing tensor contractions straightforwardly through a DSL (via CTF, tiledArray, and similar libraries), can be counterbalanced by the lack of performance arising from the continuous assembly of the distributed output subtensors. In the remainder of this work, we will evaluate the performance of the dTVC algorithm without reconstruction. As previously discussed, this operation is irrelevant in the dHOPM3 because it naturally exploits the distributive property of the tensor contraction.
@@ 5.3 Distributed native tensor–vector contraction performance
Table 2 reports on the results of the dTVC ignoring the global construction of the final tensor, which allow us to focus exclusively on the computational kernel performance. In particular, we are concerned about the matriziced tensor’s shape —tall and skinny vs. short and fat matrices— on the contraction performance. In addition to this, we propose the comparison of the looped approach based on state-of-the-art MKL kernels against our CPU-native implementation using two compilers: icpx (Intel) and clang++ (LLVM). We remind the reader that results are averaged over all contraction modes and all splitting dimensions and normalized by the theoretical peak bandwidth corresponding to eight MPI processes on MN4 (order 7–9 tensors exploit lower MPI processes due to the one-dimensional optimal splitting, as previously discussed).
It can be readily seen on Table 2 that the looped dTVC algorithm based on MKL’s batch-strided kernel yields the worst results, with relatively low average bandwidth and high standard deviation rates. The high variability of the results, except in the matrix case, evidence the lack of mode-oblivious properties of this implementation despite the fact that it is based upon a carefully optimized kernel. As demonstrated in Ref. [6], MKL is more sensitive to the shape of the matricized tensor and, in the case of concatenated small MVMs, LIBXSMM may be able to report slightly better figures. In any case, they are not expected to remain competitive against those shown in the next two columns, which belong to our CPU-native dTVC implementation using the same compilation stack as well as the LLVM infrastructure. Differences between compilers are minimum and, in both cases, our proposed algorithm achieves almost three quarters of the theoretical peak performance (1024 GB/s) with a standard deviation below 10%. We remind that this low value accounts for changes not only in the contraction mode, but also in the splitting dimension which ultimately affects the shape of the matricized form of the tensors. Therefore, we can affirm that the distributed strategy proposed in this work remains oblivious to these two variables.
Table 2: dTVC normalized, averaged bandwidth (measured as a percentage of the theoretical peak value) and corresponding unbiased, sample standard deviation percentage (within brackets) over all contraction modes and all splitting dimensions for the tensors of Table 1. Results correspond to 8 MPI processes on MN4 and two TVC algorithms (looped and native).
Order Looped (MKL) Native (icpx) Native (clang++)
@@
Figure 5(a) shows that, in general, the averaged bandwidth of all tensor contractions is between 70% to 85% of the theoretical peak on MN4, which is in good agreement with the STREAM benchmark on a single NUMA node of this architecture (81.6%). This rather small performance variability of about 15% is inevitable: for instance, not all the tensors have the exact same memory footprint as reported in Table 1. With more than 16 MPI processes, the strong scalability performance begins to increase slightly as the bandwidth bottleneck moves from main memory to L3 cache. Figure 5(b) shows absolute dTVC performance values in the form of an averaged throughput measured in kernel invocations per second (it/s). We can see how the curves associated with each tensor are very close to each other except for $d=9,10$ due to their larger memory footprint (in relative terms: 6.7% and 22.7%, respectively). Note also that these two higher-order tensors cannot be computed with more than 10 and 8 MPI processes, respectively, when using one-dimensional splitting. We confirm the linear (and superlinear) scalability already predicted by the bandwidth curves. In the light of the results gathered in Figs. 5(a)–(b), it is the first time that a distributed tensor–vector contraction remains oblivious to MPI splitting, contraction mode, and also tensor order, provided that the tensor size does not vary significantly.
@@
Figures 5(c)–(f) report on the dTVC performance for the CTE-ARM and CTE-POWER systems, respectively. We use the CPU-native, TVC algorithm to benchmark the ARM architecture while GPUs are evaluated against a looped TVC implementation making use of cuBLAS. Starting with the CPU case, Fig. 5(c) shows slightly decreasing values of the normalized bandwidth, especially for $d>3$, but they are overall within the range of 50% to 70% of the theoretical peak (on average, STREAM reports 60% on a single node). The almost linear throughput curves of Fig. 5(d) are a bit more spread than those of MN4 but, once more, we can say that our novel implementation remains oblivious to splitting, contraction mode and tensor order for this particular architecture. Finally, note that although CTE-ARM starts with better absolute throughput values over MN4 thanks to its high-end memory, its overall worse strong scalability yields subpar performance with 128 MPI processes.
@@
Figure 5(e) shows the dTVC performance on GPUs using state-of-the-art cuBLAS kernels. We observe a large variability going from 90% of the theoretical bandwidth peak to only 20% (on average, STREAM reports 61.8% on a single graphics card). Interestingly, only the last two higher-order tensors report somewhat constant bandwidth figures, but 10% to 15% below the reference STREAM benchmark value. This is expected since batched kernels do incur some overhead while enqueueing consecutive multiplication kernels within a unique function call. This distributed GPU implementation based on looped TVC lacks the obliviousness of its CPU counterpart implementation. In absolute terms, see Fig. 5(f), the CTE-POWER attains maximum throughput values around 4096 it/s, a quantity that is not much higher than the one reported by a general-purpose architecture endowed with DDR4 memory. It is worth highlighting that dTVC remains an embarrassingly parallel application (the final disjoint union operation is discarded) with zero MPI communication parts and, consequently, the scalability drop in Fig. 5(f) is likely to be attributed to the rather low device occupancy. This supports our choice of one-dimensional tensor splittings since very fine-grain parallelization tends to hamper performance, especially on this type of accelerators.
5.4 Distributed, native higher-order power method performance
@@
Table 3 confirms the conclusions previously drawn in Section 4.3 by demonstrating how a genuine data-flow execution based on either OpenMP tasks (OMPtk) or OmpSs-2 tasks (OSStk) can provide, on average, up to a 10% increase in memory bandwidth over a canonical strategy using fork-join, parallel for constructs (OMPfj). OpenMP tasks are slightly faster than OmpSs-2 tasks in this particular case, which can be attributed to the design differences between these two runtimes. It is worth mentioning that these moderate speedups are limited by the synchronous MPI collectives at the end of each external iteration of dHOPM3. In cases where the HOPM requires many iterations to reach converge ($i\gg 1$), task-based relaxation techniques similar to those that have been put in place in classical linear algebra iterative algorithms [33] could be employed to remove global synchronizations and approximate the base of output vectors. Such a case is beyond the scope of this work, but represents a real scenario where task-based parallelization could yield additional advantages.
Table 3: dHOPM3 normalized bandwidth (measured as a percentage of the theoretical peak value) over the last splitting dimension for the tensors of Table 1. Results correspond to 8 MPI processes on MN4 and three parallel implementations.
@@
Figure 6(a) shows the normalized bandwidth of dHOPM3 on MN4, which presents similarities with the dTVC bandwidth results previously shown in Fig. 5(a). The variability range is expected to widen due to the continuous decrease of the tensor size after each contraction but, in general, normalized values are between 60% to 85% most of the time. Contrarily to dTVC, dHOPM3 is not an embarrassingly parallel application and, sometimes, the cost of synchronous communications can be as important as computations as it is the case with $d=2$ and $p>32$. When looking at the absolute throughput, see Fig. 6(b), it starts at about 4–8 it/s and reaches between 512–1024 it/s with 128 MPI processes for $d=3,4$ while it drops by a factor of two for the matrix case due to excessive synchronization. It can be readily seen that Figs. 5(b) and 6(b) are reasonably equivalent with the particularity that there exists a factor above two between them. This is because the computational time of dHOPM3 is primarily dominated by two contractions, i.e. $k=1$ for $j=0$ and $k=0$ for $j=1$, as illustrated in Fig. 3(a).
@@
Figures 6(c)–(f) show the dHOPM3 results on the other two architectures equipped with high bandwidth memory. In the case of the CTE-ARM cluster, each contraction makes use of our CPU-native TVC algorithm and no significant differences can be appreciated between Figs. 5(c) and 6(c). Nevertheless, it is true that, in the case of lower-order tensors, dHOPM3 bandwidth values are more sensible to the number of processes due to synchronous communication. A similar factor above 2×2$\times$2 × between dTVC and dHOPM3 throughputs is found on CTE-ARM. Looking at the results obtained on these two CPU-based architectures, it can be inferred that the dHOPM3 algorithm inherits the obliviousness of the native dTVC kernel it is built upon.
@@
Figure 6(e) shows the dHOPM3 normalized bandwidth under Volta GPUs. Compared to the dTVC results of Fig. 5(e), dHOPM3 performance is better in the sense that curves do remain closer to each other, especially with moderate values of $p$. This is in contrast to what is observed on the two previous architectures. As a reminder, dTVC results were averaged over all the splitting dimensions while dHOPM3 figures are only calculated for the last splitting dimension for minimal memory footprint. For this reason, one can expect higher bandwidth values for dHOPM3 up to a certain number of processes; otherwise, the relatively low GPU occupancy combined with the increasing synchronization costs of collectives results in an important reduction of bandwidth, with values as low as 10% with 128 MPI processes. It is worth noting that these are indeed the best results one can achieve in terms of communication: memory is transferred between devices directly via the NCCL library that provides genuine nonblocking calls to MPI functions from the CPU host. Other communication approaches, not shown here for the sake of brevity, always resulted in subpar performance. Finally, Fig. 6(f) confirms that the maximum throughput achieved on GPUs is above 512 it/s and does not necessarily correspond to the largest number of processes. Considering all the strong scalability results presented in this section, none of the clusters equipped with high bandwidth memory is able to surpass a general purpose supercomputer with regular DDR4 memory. This is mainly attributed to the low-capacity HBM2 modules installed on these systems.
@@
Figure 7: dHOPM3 normalized throughput for the tensors of Table 1 stored in different floating-point precision formats and split along their last dimension. Results correspond to 8 MPI processes on (a) MN4 and (b) CTE-ARM.